# Assessment of the Earth orientation parameter accuracy from concurrent VLBI observations

### v1

@@ 6.4 Formal and informal errors: what did we learn?
It was long known that atmospheric turbulence cannot be described as an uncorrelated random process [see, for in instance, 36]. However, a quantitative assessment of the impact of an incorrect stochastic model used in data analysis requires significant resources for acquiring redundant data and their analysis.
@@
As we see, the use of the additive weight update did not solve the problem of underestimation of $E_{3}$ angles following the error propagation law, but only alleviated the problem. Luati and Proietti [12] proved a theorem stating that least squares with a full a priori covariance matrix is equivalent to least squares with a diagonal weight matrix if and only if certain conditions are met, and weights and observation equations that we use in our solutions do not satisfy these conditions.
@@ 6.5 Observations of different durations
If the dependence of EOP accuracy on session duration were described by the power law -0.5, the distribution of long and short VLBI observing blocks over a given interval would not have mattered The broken power law of the dependence of EOP accuracy implies that planning observing sessions longer than 2–4 hours we get a diminishing return on used resources. Extending observing sessions from 4 to 24 hours, we have to spend 6 times more resources, but we gain only 70% in accuracy instead of 140%, because the atmospheric turbulence added a correlated noise. Adding more stations does not improve noticeably the accuracy of $E_{3}$ determination.
@@ 6.6 Comparison with IERS C04 time series
It is customary to compare EOP estimates against IERS C04 series [2]. Therefore, it is instructive to compare our estimates of EOP accuracy from analysis of concurrent VLBI experiments against the C04 time series. The time series of $E_{1}$ and $E_{2}$ angles in C04 time series are dominated by the IGS final data product from processing independent GNSS phase and pseudorange data. Therefore, one can expect that the statistics of the rms of differences against C04 will be close to our estimates of accuracy based on the three corner method. Indeed, it turned out very close for vo campaigns, but 13–17% higher for r1r4 campaigns. Time series $E_{3}$ in C04 are derived from the results of processing the same VLBI group delays. According to Bizouard et al. [2], the time series of $E_{3}$ angles are derived from a) $E_{3}$ estimates from analysis of R1 and R4 VLBI experiments by 10 analysis centers and then averaged with some weights by the so-called International VLBI Service for Geodesy and Astrometry (IVS) combination center; b) $E_{3}$ estimates from some 1-hr programs: 4×4$\times$4 × daily from the Institute of Applied Astronomy in Sankt-Peterburg, Russia, and several programs processed by the United States Naval Observatory (USNO) on days when there is no r1r4 experiment. The input time series are then smoothed by using the low-pass filter. Therefore, the IERS C04 accumulates errors due to deficiencies in data analysis of individual groups and suppresses, to some degree, fast variability in EOPs due to smoothing.
@@
$E_{2}$ vo 24 0.57 0.67
$E_{1}$ r1r4 24 0.34 0.34
$E_{2}$ r1r4 24 0.32 0.35
$E_{3}$ vo/r1r4 24 0.44 0.74
$E_{3}$ K2Ws 1 1.30 1.52
$E_{3}$ MkWz 1 1.21 1.60
$E_{3}$ KkWz 1 1.50 1.48
It turned out that the differences of $E_{3}$ angles from 24-hr experiments are a factor of 1.7 greater than the accuracy of these angle estimates derived from our processing concurrent vo/r1r4 experiments (compare columns 3 and 4 in Table 9). In general, the accuracy of a combined series like IERS C04 is governed by the accuracy of input series and smoothing. We can rule out smoothing errors since the epochs of our EOP estimates and the EOP estimates of IVS combined series are practically the same. Therefore, we conclude that the IVS combined $E_{3}$ angle series are a factor of 1.7 less precise than our results. The differences of $E_{3}$ angles derived from 1-hr experiments with respect to the C04 series are a factor of 1.17–1.32 greater than the accuracy of these angle estimates from concurrent experiments, except for the time series from KkWz baselines.
In order to learn more, we took the IVS combined EOP series [^4] https://www.ccivs.bkg.bund.de/data/QUAT/COMBI/ivs24q3e.eops accessed on May 16, 2025.. We computed the statistics of the differences of IVS Combined EOP from 148 r1r4 experiments with respect to our solution after removal of the linear trend: the rms of the differences were 0.31, 0.38, and 0.86 nrad for $E_{1}$, $E_{2}$, and $E_{3}$ respectively. Similar statistics for EOP derived from 148 vo experiments are 0.61, 0.66, and 0.99 nrad.
To dig even deeper, we compared EOP estimates from the same list of concurrent vo and r1r4 observing programs using the IVS combined time series. The epochs of EOP in the IVS combined time series are close, within 2 hours, but not the same. We propagated the EOP on the common epoch using provided estimates of EOP rates. The rms of the differences divided by $\sqrt{2}$ are shown in Table 10. We see that our solution provides the rms of the differences a factor of 1.3 smaller for the $E_{3}$ angle. That prompts us to interpret the relatively large differences in IVS Combined EOP with respect to our solution as a deficiency of the IVS Combined series.
@@
nrad nrad
$E_{1}$ 0.52 0.68
$E_{2}$ 0.71 0.69
$E_{3}$ 0.44 0.57
6.7 IVS Combined EOP time series and IERS C04
We established earlier that the IVS Combined EOP series has a notable deficiency. Since the time series from r1r4 experiments are used for IERS C04, it inherits this deficiency. We should note that the IVS combined solutions fall into the category of operational solutions, while our solutions fall into the category of re-processing. We base our analysis on 50 years of efforts to improve the VLBI data analysis technique pursued by NASA Goddard Space flight center, and our data analysis technique differs from the technique commonly used by other groups for operational data analysis. In particular, we see no biases, no systematic patterns when we compare EOP from 24-hr and 1-hr experiments or EOP from different observing sessions.
Computing a combined solution, i.e. the weighted mean of results, may have some advantage for operational time series because it helps to detect a failure in quality control of an individual analysis center and allows us to mitigate situations when a given analysis center is offline. On the other hand, processing the same dataset several times and then averaging results contradicts Gauss-Markov theorem, and that causes a deviation of such a result from the optimum.
We should bear in mind worse accuracy of a combined data series, treat it as a temporary, coarse solution, and avoid using it for science. IERS C04 time series falls into the same category of operational data products since it is based on operational solutions. We have demonstrated that EOP estimates from a uniform reprocessing using advanced theoretical model and parameter estimation technique of all accumulated data are superior. In a similar way, campaigns of reprocessing GNSS observations with an improved model [see, for instance, 3] also provide superior results.
6.8 Similar comparisons made by others
An idea of evaluation of $E_{3}$ accuracy from 24-hr and 1-hr VLBI experiments is not new. A commonly used approach is to compare $E_{3}$ against IERS C04 or similar combined series. For instance, Haas et al. [9] compared $E_{3}$ derived from twelve 1-hr experiments spanned over three months against the USNO Bulletin B. They presented rms of the differences from 11 solutions in a 1.7–2.3 nrad range (their Table 2). Schartner et al. [34] made a comparison of $E_{3}$ angle estimates from several 1 hr Intensive experiments are reported rms differences against JPL EOP2 and IERS C04 time series (their Tables 1 and 2) in a range of 1.3–5.6 nrad. Wang and Xu [42] computed the rms of differences in $E_{3}$ angles determined from 1-hr observing sessions at KkWz baseline and IERS C04 experiments over 2000–2011: 1.86 nrad. Yao et al. [44] compared $E_{3}$ derived from 61 1-hr observing sessions at baselines KkWz and K2Ws in 2021. They made comparison of $E_{3}$ angles with respect to IERS C04 and reported rms of the differences: 2.05 and 1.83 nrad. In addition, they split 1-hr K2Ws into two 0.5 hr long blocks in the beginning and the end of an 1-hr observing session and processed them individually. Yao et al. [44] report that the rms of differences with respect to IERS C04 increased by 7% when 0.5-hr blocks at the beginning of the observing session were processed and decreased by 10% when 30 minute blocks at the end of the observing sessions were processed. According to the F-test, these differences are insignificant at the 0.1 confidence level. That means that throwing away 1/2 of K2Ws data did not change results. This finding contradicts to the broken power law that we derived by thinning 24-hr datasets. Inspired by these results, we performed our analysis. Our analysis showed that throwing away half of the data degraded accuracy, but only at 9% instead of 41% if the noise were uncorrelated.
Raut et al. [23] compared $E_{3}$ angle estimates determined from processing 1-hr and 24-hr experiments at three independent campaigns during 15 days in November/December 2017. They performed a number of solutions, including splitting 24-hr experiments into 1-hr blocks that they called pseudo-intensives. They report in Table 5 the rms of differences with respect to IERS C04 1.97 nrad for 24-hr blocks and 1.85 nrad for 1-hr blocks. This finding also does not support the broken power law we found, but the rms of the differences are substantially greater than in our results.
For completeness, we mention that Gipson and Baver [7] published results of a campaign of 9 experiments with a design similar to our campaign s22. According to the campaign description, each experiment contained twenty two 1-hr scheduling blocks at KkWz baseline among with data at other stations. However, database files in the public archive at NASA Crustal Dynamics Data Information System (CDDIS) contains data at all stations, except stations kokee and wettzell. Apparently, Gipson and Baver [7] based their analysis on additional datasets that were not publicly released. Considering governing open science policies, we defer comments on these results till missing datasets will become available.
MacMillan [13] presented wrms of the differences of EOP from two concurrent VLBI networks during 15 consecutive days during CONT17 observing campaign in November/December 2017. He reports the rms of the differences in Euler angle vector (0.21, 0.19, 0.25) nrad. This is a factor of 2–3 smaller than the rms of the differences in EOP derived from analysis of group delays from vo/r1r4 experiments presented in Table 4. There are three factors that played the role. First, CONT17 experiments are not representative. This campaign was an attempt to get the best results using existing VLBI hardware, while vo and r1r4 experiments have been running as regular, operational observing sessions. Second, these experiments ran in winter time for most of the stations in the network. And third, statistics over a continuous 15 day period are different from statistics over a 5 year period in a presence of red noise. Ray et al. [25] argues strongly that the longer the time period for error characterization, the greater the errors, and we find their argumentation quite convincing. Petrov and Kovalev [22] presented strong evidence of the impact of red noise on results of comparisons of two subsets of VLBI data on source positions. At present, we are not in a position to deliver a quantitative proof of the presence of red noise in EOP estimates at scales of 0.01–5 years, but we consider that the first two factors are not sufficient to explain the discrepancies in statistics of that magnitude.
@@ 7 Summary and Outlook
Having processed a number of recent and historical datasets of concurrent VLBI observations at different networks, we have arrived at more reliable estimates of the accuracy of time variable Euler angles describing the Earth’s rotation. These estimates have a little resemblance to formal computed using the law of error propagation under the assumption that the noise is uncorrelated. We summarize the accuracy of determination of $E_{3}$ angles from various 24-hr and 1-hr campaigns in Figure 11.
@@
We found no evidence that advanced scheduling strategies for improved determination of the atmospheric path delay in zenith direction have any impact on the accuracy of EOP estimates. Inability to predict the accuracy based on simulation suggests that our approaches to data analysis and simulation of geodetic VLBI have to be revised. We recognize that this inability poses a serious problem. The use of a weight matrix with filled diagonal terms and zero off-diagonal terms was a necessity in the 19th and 20th centuries due to the scarcity of computing resources, but the use of this technique by inertia in the 21st century is an anachronism. We are going to enhance our data analysis procedure by the use of a full weight matrix, including off-diagonal terms, and encourage others to pursue these efforts. The goal of these efforts is to predict accuracy of EOP estimates within tens of percent, explain the power law -0.3 and seasonality in EOP accuracy, and explain correlation 0.46 between estimates of $E_{3}$ angles from 0.5-hr data segments. The dataset presented in this study is an excellent testbed that can be used for tuning algorithms that eventually are supposed to produce correct estimates of uncertainties of estimated parameters in the presence of a correlated noise.
@@ Acknowledgements
We would like to thank Frank Lemoine for valuable comments. It is our pleasure to thank staff at the CDDIS for making available Level 2 VLBI data. We thank Cynthia Thomas and Dirk Behrend for coordinating International VLBI for Geodesy and Astrometry observing programs.
@@
  • Code availability: we used for our analysis open source software package SGDASS available at https://doi.org/10.25966/7z5f-eg07. The s22 schedules are generated using the VieSched++ software package [31] available at https://github.com/TUW-VieVS/VieSchedpp under the GNU General Public License.

### v2

@@ 6.5 Formal and informal errors: what did we learn?
It was long known that atmospheric turbulence cannot be described as an uncorrelated random process [see, for in instance, r:tat71]. However, a quantitative assessment of the impact of an incorrect stochastic model used in data analysis requires significant resources for acquiring redundant data and their analysis.
@@
In order to see how the use of additive weights affected EOP accuracy, we ran a solution without weight updates. The accuracies increased by factors of 1.42, 1.23, and 1.41 for $E_{1}$, $E_{2}$, and $E_{3}$ components respectively. As we see, the use of the additive weight update did not solve the problem of underestimation of $E_{3}$ angles following the error propagation law, but only alleviated the problem.   r:lua08 have proven a theorem that established the necessary and and sufficient conditions of equivalence of a least square solution with a diagonal weight matrix with a solution with a full a priori covariance matrix: if there are $p$ linear combinations of columns of observation matrix that are eigenvector of $(\mathop{\rm Cov}\nolimits(\epsilon,{\epsilon}^{{}^{\bf\top}}\!)W)$, where $p$ is the number of parameters, $(\mathop{\rm Cov}\nolimits(\epsilon,{\epsilon}^{{}^{\bf\top}}\!)$ is the covariance matrix of right-hand sides and $W$ is the diagonal weight matrix. It follows from the theorem that an update of diagonal elements is not sufficient to satisfy condition.
@@ 6.6 Observations of different durations
If the dependence of EOP accuracy on session duration were described by the power law -0.5, the distribution of long and short VLBI observing blocks over a given interval would not have mattered The broken power law of the dependence of EOP accuracy implies that planning observing sessions longer than 2–4 hours we get a diminishing return on used resources. Extending observing sessions from 4 to 24 hours, we have to spend 6 times more resources, but we gain only 70% in accuracy instead of 140%, because the atmospheric turbulence added a correlated noise.
@@
6.7 Similar comparisons made by others
An idea of evaluation of $E_{3}$ accuracy from 24-hr and 1-hr VLBI experiments is not new. A commonly used approach is to compare $E_{3}$ against IERS C04 or similar combined series. For instance, r:haa21 compared $E_{3}$ derived from twelve 1-hr experiments spanned over three months against the USNO Bulletin B. They presented rms of the differences from 11 solutions in a 1.7–2.3 nrad range (their Table 2). r:scha22 made a comparison of $E_{3}$ angle estimates from several 1 hr Intensive experiments are reported rms differences against JPL EOP2 and IERS C04 time series (their Tables 1 and 2) in a range of 1.3–5.6 nrad. r:wan12 computed the rms of differences in $E_{3}$ angles determined from 1-hr observing sessions at KkWz baseline and IERS C04 experiments over 2000–2011: 1.86 nrad. r:yao23 compared $E_{3}$ derived from 61 1-hr observing sessions at baselines KkWz and K2Ws in 2021. They made comparison of $E_{3}$ angles with respect to IERS C04 and reported rms of the differences: 2.05 and 1.83 nrad. In addition, they split 1-hr K2Ws into two 0.5 hr long blocks in the beginning and the end of an 1-hr observing session and processed them individually. r:yao23 report that the rms of differences with respect to IERS C04 increased by 7% when 0.5-hr blocks at the beginning of the observing session were processed and decreased by 10% when 30 minute blocks at the end of the observing sessions were processed. According to the F-test, these differences are insignificant at the 0.1 confidence level. That means that throwing away 1/2 of K2Ws data did not change results. This finding contradicts to the broken power law that we derived by thinning 24-hr datasets. Inspired by these results, we performed our analysis. Our analysis showed that throwing away half of the data degraded accuracy, but only at 9% instead of 41% if the noise were uncorrelated.
r:raut22 compared $E_{3}$ angle estimates determined from processing 1-hr and 24-hr experiments at three independent campaigns during 15 days in November/December 2017. They performed a number of solutions, including splitting 24-hr experiments into 1-hr blocks that they called pseudo-intensives. They report in Table 5 the rms of differences with respect to IERS C04 1.97 nrad for 24-hr blocks and 1.85 nrad for 1-hr blocks. This finding also does not support the broken power law we found, but the rms of the differences are substantially greater than in our results.
For completeness, we mention that r:gip16 published results of a campaign of 9 experiments with a design similar to our campaign s22. According to the campaign description, each experiment contained twenty two 1-hr scheduling blocks at KkWz baseline among with data at other stations. However, database files in the public archive at NASA Crustal Dynamics Data Information System (CDDIS) contains data at all stations, except stations kokee and wettzell. Apparently, r:gip16 based their analysis on additional datasets that were not publicly released. Considering governing open science policies, we defer comments on these results till missing datasets will become available.
r:mcm22 presented wrms of the differences of EOP from two concurrent VLBI networks during 15 consecutive days during CONT17 observing campaign in November/December 2017. He reports the rms of the differences in Euler angle vector (0.21, 0.19, 0.25) nrad. This is a factor of 2–3 smaller than the rms of the differences in EOP derived from analysis of group delays from vo/r1r4 experiments presented in Table 4. There are three factors that played the role. First, CONT17 experiments are not representative. This campaign was an attempt to get the best results using existing VLBI hardware, while vo and r1r4 experiments have been running as regular, operational observing sessions. Second, these experiments ran in winter time for most of the stations in the network. And third, statistics over a continuous 15 day period are different from statistics over a 5 year period in a presence of red noise. r:ray17 argues strongly that the longer the time period for error characterization, the greater the errors, and we find their argumentation quite convincing. r:rfc1 presented strong evidence of the impact of red noise on results of comparisons of two subsets of VLBI data on source positions. At present, we are not in a position to deliver a quantitative proof of the presence of red noise in EOP estimates at scales of 0.05–5 years, but we consider that the first two factors are not sufficient to explain the discrepancies in statistics of that magnitude.
@@ 6.8 Comparison with IERS C04 time series
As we see from the previous section, it is customary to compare EOP estimates against IERS C04 series [r:iers_c04], implicitely assuming that C04 represents the groud truth, and therefore, the differences against C04 can be considered as a measure of accuracy.   Since the reported statsitics of the differences with respect to IERS C04 disagree with our estimates of EOP accuracy based on processing concurrent observations, it is instructive to compare our   EOP estimates against the C04 time series. The time series of $E_{1}$ and $E_{2}$ angles in C04 time series are dominated by the IGS final data product from processing independent GNSS phase and pseudorange data. Therefore, one can expect that the statistics of the rms of differences against C04 will be close to our estimates of accuracy based on the three corner method. Indeed, it turned out very close for vo campaigns, but 13–17% higher for r1r4 campaigns. Time series $E_{3}$ in C04 are derived from the results of processing the same VLBI group delays. According to r:iers_c04, the time series of $E_{3}$ angles are derived from a) $E_{3}$ estimates from analysis of R1 and R4 VLBI experiments by 10 analysis centers and then averaged with some weights by the so-called International VLBI Service for Geodesy and Astrometry (IVS) combination center; b) $E_{3}$ estimates from some 1-hr programs: $4\times$ daily from the Institute of Applied Astronomy in Sankt-Peterburg, Russia, and several programs processed by the United States Naval Observatory (USNO) on days when there is no r1r4 experiment. The input time series are then smoothed by using the low-pass filter. Therefore, the IERS C04 accumulates errors due to deficiencies in data analysis of individual groups and suppresses, to some degree, fast variability in EOPs due to smoothing.
@@
$E_{2}$ vo 24 0.57   0.72
$E_{1}$ r1r4 24 0.34   0.39
$E_{2}$ r1r4 24 0.32   0.42
$E_{3}$ vo/r1r4 24 0.40   0.75
$E_{3}$ K2Ws 1   1.40 1.52
$E_{3}$ MkWz 1   1.32 1.60
$E_{3}$ KkWz 1   1.59 1.48
It turned out that the differences of $E_{3}$ angles from 24-hr experiments are a factor of   1.8 greater than the accuracy of these angle estimates derived from our processing concurrent vo/r1r4 experiments (compare columns 3 and 4 in Table 9). In general, the accuracy of a combined series like IERS C04 is governed by the accuracy of input series and smoothing.   Smoothing is not expected to be a major factor, since the epochs of our EOP estimates and the EOP estimates of IVS combined series are practically the same. Therefore, we conclude that the IVS combined $E_{3}$ angle series are a factor of 1.8 less precise than our results. The differences of $E_{3}$ angles derived from 1-hr experiments with respect to the C04 series are a factor of   1.1–1.2 greater than the accuracy of these angle estimates from concurrent experiments, except for the time series from KkWz baselines.
We have a disagreement with comparisons made by others. In order to understand the origin of this disagreement, we took the IVS combined EOP series [^4] https://www.ccivs.bkg.bund.de/data/QUAT/COMBI/ivs24q3e.eops accessed on May 16, 2025. and computed the statistics of the differences of IVS Combined EOP from 148 r1r4 experiments with respect to our solution after removal of the linear trend: the   standard deviations of the differences were 0.31, 0.38, and 0.86 nrad for $E_{1}$, $E_{2}$, and $E_{3}$ respectively. Similar statistics for EOP derived from 148 vo experiments are 0.61, 0.66, and 0.99 nrad. Statistics of the standard deviations in differences between IVS and IERS time series over 265 vo/r1r4 experiments are 0.56, 0.62, and 1.09 nrad.
To dig even deeper, we compared EOP estimates from the same list of concurrent vo and r1r4 observing programs using the IVS combined time series. The epochs of EOP in the IVS combined time series are close, within 2 hours, but not the same. We propagated the EOP on the common epoch using provided estimates of EOP rates. The rms of the differences divided by $\sqrt{2}$ are shown in Table 10. We see that our solution provides the rms of the differences a factor of 1.4 smaller for the $E_{3}$ angle.   Therefore, we conclude that the relatively large differences in IVS Combined EOP time series with respect to our solution as a deficiency of the IVS Combined series.
@@
bias rms bias rms
nrad nrad nrad nrad
$E_{1}$ 0.20   0.54 0.60 0.68
$E_{2}$ 0.51 0.71 -0.44 0.69
$E_{3}$ 0.14   0.40 0.51 0.57
@@ 7 Summary and Outlook
Having processed a number of recent and historical datasets of concurrent VLBI observations at different networks, we have arrived at more reliable estimates of the accuracy of time variable Euler angles describing the Earth’s rotation. These estimates have a little resemblance to formal computed using the law of error propagation under the assumption that the noise is uncorrelated. They also noticeably smaller than the differences against IERS C04 time series that was often used as a measure of EOP accuracy in literature. We summarize the accuracy of determination of $E_{3}$ angles from various 24-hr and 1-hr campaigns in Figure 10.
@@
Based on statistics of differences in positions of a large sample of sources determined from analysios of observations at different bands reported in literature, we got the quantative estimate of the upper limit of the impact of source structure on EOP estimates: 0.1 nrad. Comparing the extra variance in EOP estimates in summer versus in winter with the impact of source structure on EOP, we conclude that the impact of source structure contribution is one order of magnitude less than the impact of unmodeled atmospheric path delay.
We found no evidence that advanced scheduling strategies for improved determination of the atmospheric path delay in zenith direction have any impact on the accuracy of EOP estimates. Inability to predict the accuracy based on simulation suggests that our approaches to data analysis and simulation of geodetic VLBI have to be revised. We recognize that this inability poses a serious problem. r:tiege25 simulated an increase of repeatabilities in EOP estimates when 1/4 observations is removed, and they found that $E_{3}$ errors defined as repeatabilities increased by 18.8%. In contrast, our analysis of real data showed that $E_{3}$ accuracy from 18 h experiments differs by less than 1% with respect to the accuracy derived from analysis of 24 hr experiments.
The use of a weight matrix with filled diagonal terms and zero off-diagonal terms was a necessity in the 19th and 20th centuries due to the scarcity of computing resources, but the use of this technique by inertia in the 21st century is an anachronism. We are going to enhance our data analysis procedure by the use of a full weight matrix, including off-diagonal terms, and encourage others to pursue these efforts. The goal of these efforts is to predict accuracy of EOP estimates within tens of percent, explain the power law -0.3 and seasonality in EOP accuracy, and explain correlation 0.46 between estimates of $E_{3}$ angles from 0.5-hr data segments. The dataset presented in this study is an excellent testbed that can be used for tuning algorithms that eventually are supposed to produce correct estimates of uncertainties of estimated parameters in the presence of a correlated noise.
@@ Acknowledgements
We would like to thank Frank Lemoine, Christian Bizouard, and Chris Jacobs for valuable comments. It is our pleasure to thank staff at the CDDIS for making available Level 2 VLBI data. We thank Cynthia Thomas and Dirk Behrend for coordinating International VLBI for Geodesy and Astrometry observing programs. We also would like to express our gratitude to the staff of observing geodetic stations, VLBI correlators, technology development groups, and data centers working together and forming IVS for their dedication. The overview of IVS can be found in [r:ivs].
@@
  • Code availability: we used for our analysis open source software package SGDASS available at https://doi.org/10.25966/7z5f-eg07. The s22 schedules are generated using the VieSched++ software package [r:scha19] available at https://github.com/TUW-VieVS/VieSchedpp under the GNU General Public License.
@@
Appendix A Scheduling strategies
While generally speaking, observations at longer baselines are more sensitive for measuring Earth’s rotation angle $E_{3}$, they also face limitations due to a reduced common visibility of the sky [r:scha21]. The $\sim\!8400$ km baseline between stations Mg and Ws presents certain geometric advantages. At this distance, observations near the zenith at one station correspond to observations at very low elevation angles at the other station (see Figure 11). This geometry is particularly beneficial for solving for residual atmospheric path delay in zenith direction, which remains the primary source of error in VLBI measurements. To effectively separate atmospheric zenith path delay from other parameters, such as station height and clock biases, it is essential to observe at a wide range of elevation angles.
Figure 11: Mutual visibility between Mg (left) and Ws (right) stations. The elevation of the other station is color-coded. The black line marks the visible horizon [r:scha23].
The s22 sessions were specifically designed to investigate improved estimation of residual atmospheric path delay in zenith direction. Each session is divided into 1-hour blocks that alternate between two scheduling strategies. Strategy A follows the conventional scheduling approach utilized in the VieSched++ scheduling software [r:scha19] and serves as a baseline reference. By contrast, strategy B is tailored to maximize the variability in elevation angles, promoting frequent alternation between high- and low-elevation observations to improve sensitivity to zenith atmospheric path delay parameters.
To implement Strategy B, sources at high elevation for each station are identified in advance. Scans are then scheduled according to the following repeating pattern:
  1. 1.
    a source at high elevation for Mg (corresponds to low elevation for Ws),
  2. 2.
    any source,
  3. 3.
    a source at high elevation for Ws (corresponds to low elevation for Mg),
  4. 4.
    any source.
The interleaved “any source” steps (2 and 4) help to maintain a balanced distribution of observations across different sources. In general, a minimum interval of 20 minutes is enforced between consecutive observations of the same source. However, due to the limited availability of high-elevation sources at each station, this restriction is relaxed to 10 minutes for those specific sources in strategy B. Additionally, to ensure adequate coverage of low-declination regions, a low-declination source is explicitly scheduled every 15 minutes in both strategies. The impact of this strategy on the distribution of observation is discussed in [r:scha23].
Appendix B Relationship between daily estimates of EOP in different notations
We used in our analysis a simple mathematical model in expression 13 to describe dependence of time of residual Euler angles with Earth Orientation Parameters $P_{c},P_{s},S_{c},S_{s},R_{c},R_{s},E_{i}(s),\dot{E}_{i}(s)$. An alternative mathematical model was introduced by [r:her86a]:
$\vec{q}_{e}(t)=\left(\begin{array}[]{llll}\mu(\,Y_{p}(s)&+\dot{Y_{p}}(s)\cdot(t-t_{\rm ref})&\hskip-65.00009pt+\Delta X(s)\cos-\Omega_{n}t&+\Delta Y(s)\sin-\Omega_{n}t)\vskip 2.15277pt\\ \mu(\,X_{p}(t)&+\dot{X_{p}}(s)\cdot(t-t_{\rm ref})&\hskip-65.00009pt+\Delta X(s)\sin-\Omega_{n}t&-\Delta Y(s)\cos-\Omega_{n}t)\vskip 2.15277pt\\ \kappa(\,{\rm UT1-UTC}(s))&+\frac{\displaystyle d}{\displaystyle dt}({\rm UT1-UTC})(s)\cdot(t-t_{\rm ref}))&&\\ \end{array}\right),$ (27)
where $X_{p}$, $Y_{p}$, polar motion, $\Delta X$, $\Delta Y$ are pole offsets, and $\mu$ and $\kappa$ are scaling factors introduced in section 2.
Comparing with expression 13, we immediately see that
[…]
We put sign $\approx$ because the two parameterizations of harmonic variations are not equivalent. The model in expression 27 is less precise. The differences in the first order approximation are
$\displaystyle\frac{\displaystyle 1}{\displaystyle t_{b}-t_{e}}\,\int\limits_{t_{b}}^{t_{e}}\>\sum_{j}^{N}P^{s}_{j}\sin((\omega_{j}+\Omega_{n})\,t+\phi_{j})\,dt,$ (35)
where $t_{b}$ and $t_{e}$ are start and stop of an experiment. These differences are small for spectrum constituents with frequencies close to $\Omega_{n}$, i.e nutations with long periods viewed in the celestial coordinates system. Celestial pole offsets can be related to $P^{c},P^{s},S^{c},S^{s}$ as
[…]
where summation is performed only over constituents of the spectrum in the retrograde diurnal band. We should note that no $R^{s},R^{c}$ terms are accounted when parameterization in equation 27 used. If the a priori model of the dependence of Euler angles on time is known precisely, the difference between this parameterization and the parameterization in equation 13 is negligible. However, errors of the a priori model will affect $E_{1}(s),E_{2}(s),E_{3}(s)$ and $X_{p}(s),Y_{p}(s)$, and UT1-UTC(s) differently.